OpenStudy (anonymous):

mommy raised $30,000 for a sports fund. No more money will be placed in the fund. Each year the fund will decrease by 5%. Determine the amount of money, to the nearest cent, that will be left in the sports fund after 4 years.

OpenStudy (dumbcow):

B = P(1-i)^n B = 30,000(.95)^4 B = 24,435.19

OpenStudy (anonymous):

That's the formula! Thank you so much (:

OpenStudy (dumbcow):

your welcome